# Integral means spectrum functionals on Teichmüller spaces

### v8

@@ Abstract.
In this paper we introduce and study the integral means spectrum (IMS) functionals on Teichmüller spaces. We show that the IMS functionals on the closure of the universal Teichmüller space and the universal asymptotic Teichmüller space are both continuous. During the proof, we obtain some new results about the universal asymptotic Teichmüller space. We also show that the integral means spectrum of any univalent function admitting a quasiconformal extension to the extended complex plane is strictly less than the universal integral means spectrum.
@@ 1. Introduction
We say a function $f\in\mathcal{S}$ is an extremal function for B(t)B(t)(or $B_{b}(t)$) if $\beta_{f}(t)=B(t)$ (or $\beta_{f}(t)=B_{b}(t)$).
@@
We let $\mathcal{S}_{q}$ be the class of all univalent functions $f$ that belong to $\mathcal{S}_{b}$ and admit a quasiconformal extension to $\widehat{\mathbb{C}}$. By the fractal approximation principle, see [24], [8], we know that
@@
We see from Theorem 1.12 that finding out the exact values of $B_{b}(t)$ can be thought of as an extremal problem in the class $\mathcal{S}_{q}$. In this paper, we introduce and study some functionals, induced by the integral means spectrum $\beta_{f}(t)$, on the universal Teichmüller space $T$ and the universal asymptotic Teichmüller space $AT$. We call them integral means spectrum (IMS) functionals. We will show that
@@
The paper is organized as follows. We recall some basic definitions and results on univalent functions and quasiconformal mappings and then prove some auxiliary results in the next section. In Section 3, after introducing some definitions and notations, we will restate Theorem 1.13, 1.14, 1.15. We shall present the proof of Theorem 1.13, 1.14, 1.15 in Section 4. In Section 5, it will be shown that the integral means spectrum of any univalent function admitting a quasiconformal extension to $\widehat{\mathbb{C}}$ is strictly less than the universal integral means spectrum and some remarks and questions will also be given.
@@ 2. Preliminaries and auxiliary results
Here $\tau$ is a univalent function from $\Omega$ to $\Delta$. In particular, $\rho_{\Delta}(z)=1/(1-|z|^{2}),z\in\Delta.$ For an analytic function $f$ in the simply connected domain $\Omega$ with $f(w)\neq 0$ for all $w\in\Omega$, we will use $\log f$ to denote certain one single-valued branch of the logarithm of $f$. We will use $C(\cdot),C_{1}(\cdot),C_{2}(\cdot),\cdots$ to denote some positive numbers which depend only on the elements in the bracket and the numbers may be different in different places.
@@ 2.1. Univalent functions (conformal mappings)
Let $f$ be a locally univalent function in an open domain $\Omega$ of $\mathbb{C}$. The Schwarzian derivative $S_{f}$ of $f$ is defined as
$S_{f}(z):=[N_{f}(z)]^{\prime}-\frac{1}{2}[N_{f}(z)]^{2}=\bigg{[}\frac{f^{\prime\prime}(z)}{f^{\prime}(z)}\bigg{]}^{\prime}-\frac{1}{2}\left[\frac{f^{\prime\prime}(z)}{f^{\prime}(z)}\right]^{2},\,z\in\Omega.$
Here $N_{f}(z)=f^{\prime\prime}(z)/{f^{\prime}(z)}$ is called as the Pre-Schwarzian derivative of $f$. Let $g$ be another locally univalent function in $f(\Omega)$. Then we have
@@
It has been proved in [22, Page 115] that $\mathbf{S}$ is closed in $E_{2}$. The statement that $\mathbf{N}$ is closed in $E_{1}$ is also known in the literature. For the completeness of the paper, we will present a direct proof for this statement and some arguments of the proof will be used later. First, we have
@@
The if part of this claim is easy. We only prove the only if part. We assume that $N_{f_{j}}(z):=\phi(z),j=1,2$. Then
@@
We proceed to prove the statement. We suppose that there is a sequence $\{f_{n}\}_{n=1}^{\infty}$, $f_{n}\in\mathcal{U}$ and ${\Phi}\in E_{1}$ such that
$\lim\limits_{n\rightarrow\infty}\|N_{f_{n}}-\Phi\|_{E_{1}}=0.$
@@
$\Phi(z)=\lim\limits_{n\rightarrow\infty}N_{f_{n}}(z),$
for each $z\in\Delta$. We will show that there is an $f\in\mathcal{U}$ such that $N_{f}=\Phi$. Without loss of generality, since the Pre-Schwarzian derivative is affine invariant, we assume that $f_{n}\in\mathcal{S}$ for all $n$. Noting that $\{f_{n}\}$ is a normal family, then $\{f_{n}\}$ contains a subsequences (still denoted by $\{f_{n}\}$) which is locally uniformly convergence in $\Delta$. We denote by $g$ the limit of the subsequences, i.e.,
$g(z)=\lim\limits_{n\rightarrow\infty}f_{n}(z),$
@@
$N_{g}(z)=\lim\limits_{n\rightarrow\infty}N_{f_{n}}(z).$
On the other hand, for any $\phi\in E_{1}$, there is a locally univalent function $f_{\phi}$ in $\Delta$ such that
@@ 2.2. Quasiconformal mappings
On the other hand, by Lemma 2.9, we have
@@ 3. Integral means spectrum functionals on Teichmüller spaces
We say $\mu$ and $\nu$ in $\mathcal{M}({\Delta}^{*})$ are asymptotically equivalent if there exists some $\widetilde{\nu}$ such that $\widetilde{\nu}$ and $\nu$ are equivalent and $\widetilde{\nu}(z)-\mu(z)\rightarrow 0$ as $|z|\rightarrow 1^{+}$. The asymptotic equivalence of $\mu$ will be denoted by $\mu]_{AT}$. The asymptotic universal Teichmüller space $AT$ is the set of all the asymptotic equivalence classes $\mu]_{AT}$ of elements $\mu$ in $\mathcal{M}({\Delta}^{*})$. The Teichmüller distance $d_{AT}([\mu]_{AT},[\nu]_{AT})$ of two points $\mu]_{AT}$, $\nu]_{AT}$ in $AT$ is defined as
@@
We see from [26] that $|\Phi(t)|\leq 8$. Noting that $f(z,0)=z$ so that $\Phi(0)=0$, then by using Schwarz’s lemma, we obtain that $|\Phi(t)|\leq 8t\|\mu_{\mathbf{f}}\|_{\infty}$ for $t\in\mathcal{D}_{\mathbf{f}}$. On the other hand, from $f(z,1)=\mathbf{f}(z)$, we have $|N_{\mathbf{f}}(z)|\rho_{f_{\mu}(\Delta)}^{-1}(z)\leq 8\|\mu_{\mathbf{f}}\|_{\infty}$ for all $z\in f_{\mu}(\Delta)$. It follows from the fact $\mathbf{f}=f_{\nu}\circ f_{\mu}^{-1}$ that
@@
In view of Propposition 3.2 and 3.3, we can identify the universal Teichmüller space with $T_{1}$ or $T_{2}$. Let $\mathcal{S}_{q}^{\infty}$ be the class of all functions $f\in\mathcal{S}_{q}$ which have a quasiconformal extension, still denoted by $f$, to $\widehat{\mathbb{C}}$ satisfying that $f(\infty)=\infty$. We set
@@
From $\mathbf{h}=g\circ f^{-1}$ we have
@@
where $z_{r}=re^{i\arg z}$, and the integral is taken on the radial path from $z_{r}$ to $z$. On the other hand, since for all $|z|\in(r,1)$, $|P(z)(1-|z|^{2})|<\varepsilon$, then we have
@@
From [34, Proposition 8], we see that
@@ 5. Final results and remarks
is holomorphic in $\Delta$. For fixed $r\in(0,1)$, by [30, Theorem 2.4.8], we have
@@
When $t_{1}>0$, let $\beta:=\beta_{f_{\mu}}(t_{1})>0$. Let $\varepsilon<\beta$ be a positive number, which will be fixed later. Then we see from the definition of integral means spectrum of $f_{\mu}$ that there is a sequence $\{r_{n}\}_{n=1}^{\infty}$ with $r_{n}<1$ and $r_{n}\rightarrow 1$ as $n\rightarrow\infty$ and such that
@@
Then it is easy to see that $\mathcal{E}_{n}\bigcup\mathcal{F}_{n}=[0,2\pi)$ and $\mathcal{E}_{n}\bigcap\mathcal{F}_{n}=\emptyset.$ We denote
@@
Hence, from [5] or [2], [35], we see that $\mathbf{h}$ is bounded univalent in $\Omega$ and admits a quasiconformal extension (still denoted by $\mathbf{h}$) to $\widehat{\mathbb{C}}$ when $\varepsilon$ small enough. Now, we let $\varepsilon$ be small enough so that $\mathbf{h}$ has a quasiconformal extension to $\widehat{\mathbb{C}}$. We consider the mapping $\mathbf{F}(z)=\mathbf{h}\circ f_{\mu}(z),z\in\widehat{\mathbb{C}}$. Then $\mathbf{F}$ is a quasiconformal mapping from $\widehat{\mathbb{C}}$ to itself and it is bounded conformal in $\Delta$. We let $\mathbb{F}=\mathbf{F}|_{\Delta}$. Then we see that $\mathbb{F}$ belongs to $\mathcal{S}_{q}$ and for any $r\in(0,1)$,
$\int_{0}^{2\pi}|\mathbb{F}^{\prime}(re^{i\theta})|^{t}d\theta$ $=$ $\int_{0}^{2\pi}|\mathbf{h}^{\prime}\circ f_{\mu}(re^{i\theta})|^{t}|f^{\prime}_{\mu}(re^{i\theta})|^{t}d\theta$
$=$ $\int_{0}^{2\pi}|f^{\prime}_{\mu}(re^{i\theta})|^{t+t\varepsilon}d\theta.$
@@
This corollary tells us that the extremal function for $B_{b}(t)$ can not be from the class $\mathcal{S}_{q}$ when $t\neq 0$. Recall that Theorem 1.12 in the introduction, this result seems to be surprising.
@@
Corollary 5.5 tells us that, if the functional $I_{\overline{\mathcal{T}}}$ attains a maximum on $\overline{\mathcal{T}}$ for $t\neq 0$, then $B_{b}(t)$ must has an extremal function and the Pre-Schwarzian derivative of the extremal function for $B_{b}(t)$ must lie in $\partial\mathcal{T}$. In fact, for some special cases, we know that $I_{\overline{\mathcal{T}}}$ attains a maximum at certain point in $\partial{\mathcal{T}}$. For example,
@@
For each $t\neq 0$, $B_{b}(t)$ has an extremal function whose Pre-Schwarzian derivative lies in $\partial\mathcal{T}$.

### v11

@@ Abstract.
In this paper we introduce and study the integral means spectrum (IMS) functionals on Teichmüller spaces. We show that the IMS functionals on the closure of the universal Teichmüller space and the universal asymptotic Teichmüller space are both continuous. During the proof, we consider the Pre-Schwarzian derivative model of universal asymptotic Teichmüller space and establish some new results for it. We also show that the integral means spectrum of any univalent function admitting a quasiconformal extension to the extended complex plane is strictly less than the universal integral means spectrum.
@@ 1. Introduction
We say a function $f\in\mathcal{S}$ is an extremal function for $B_{b}(t)$(or $B(t)$) if $\beta_{f}(t)=B_{b}(t)$ (or $\beta_{f}(t)=B(t)$).
@@
We let $\mathcal{S}_{q}$ be the class of all univalent functions $f$ that belong to $\mathcal{S}_{b}$ and admit a quasiconformal extension to $\widehat{\mathbb{C}}$. Later, if a univalent function $f$ belongs to $\mathcal{S}_{q}$, we still use $f$ to denote its quasiconformal extension. Recently, some researchers have studied the integral means spectrums of univalent functions admitting a quasiconformal extension to $\widehat{\mathbb{C}}$, see for example, [20], [21], [22], [23], [24], [35] and [36]. By the fractal approximation principle, see [29], [8], we know that
@@
We see from Theorem 1.12 that finding out the exact values of $B_{b}(t)$ can be thought of as a global extremal problem in the class $\mathcal{S}_{q}$. In this paper, we do not seek to find a better value of $B_{b}(t)$ but instead focus on the following
Question 1.13.
(1) For fixed $t\in\mathbb{R}$ with $t\neq 0$, how does the integral means spectrum $\beta_{f}(t)$ depend on $f$ in the class $\mathcal{S}_{q}$? (2) For each $t\neq 0$, does there exist at least one extremal function for $B_{b}(t)$? (3) If the extremal functions for $B_{b}(t)$ exist, which subset of $S$ do they belong to?
To answer this question, we introduce and study some functionals, induced by the integral means spectrum $\beta_{f}(t)$, on the universal Teichmüller space $T$ and the universal asymptotic Teichmüller space $AT$. We call them integral means spectrum (IMS) functionals. We will show that
@@
The paper is organized as follows. We recall some basic definitions and results on univalent functions and quasiconformal mappings, and then prove some auxiliary results in the next section. In Section 3, after introducing some definitions and notations, we will restate Theorem 1.14, 1.15, 1.16. We shall present the proof of Theorem 1.14, 1.15, 1.16 in Section 4. In Section 5, we establish a final main theorem, which shows that the integral means spectrum of any univalent function admitting a quasiconformal extension to $\widehat{\mathbb{C}}$ is strictly less than the universal integral means spectrum. We also give some remarks and questions in this last section.
@@ 2. Preliminaries and auxiliary results
Here $\tau$ is a univalent function from $\Omega$ to $\Delta$. In particular,
$\rho_{\Delta}(z)=(1-|z|^{2})^{-1},z\in\Delta.$
For an analytic function $f$ in the simply connected domain $\Omega$ with $f(w)\neq 0$ for all $w\in\Omega$, we will use $\log f$ to denote a certain single-valued branch of the logarithm of $f$. We will use $C(\cdot),C_{1}(\cdot),C_{2}(\cdot),\cdots$ to denote some positive numbers which depend only on the elements in the bracket and the numbers may be different in different places.
@@ 2.1. Univalent functions (conformal mappings)
Let $f$ be a locally univalent function in an open domain $\Omega$ of $\mathbb{C}$. The Pre-Schwarzian derivative $N_{f}$ of $f$, and the Schwarzian derivative $S_{f}$ of $f$ are defined as
$N_{f}(z):=\frac{f^{\prime\prime}(z)}{{f^{\prime}(z)}},\,z\in\Omega,$
and
$S_{f}(z):=[N_{f}(z)]^{\prime}-\frac{1}{2}[N_{f}(z)]^{2}=\frac{f^{\prime\prime\prime}(z)}{f^{\prime}(z)}-\frac{3}{2}\left[\frac{f^{\prime\prime}(z)}{f^{\prime}(z)}\right]^{2},\,z\in\Omega.$
Let $g$ be another locally univalent function in $f(\Omega)$. Then we have
(2.1) $N_{g\circ f}(z)=N_{g}(f(z))[f^{\prime}(z)]+N_{f}(z),\,\,z\in\Omega,$
and
@@
It has been proved in [27, Page 115] that $\mathbf{S}$ is closed in $E_{2}$. The statement that $\mathbf{N}$ is closed in $E_{1}$ is also known in the literature. For the completeness of the paper, we will present a direct proof for this statement and some arguments of the proof will be used later. First, we have
@@
The if part of this claim is easy. We only prove the only if part. We let $N_{f_{j}}(z):=\phi(z),j=1,2$. Then
@@
We proceed to prove the statement. We suppose that there is a sequence $\{{\bf f}_{n}\}_{n=1}^{\infty}$, ${\bf f}_{n}\in\mathcal{U}$ and ${\Phi}\in E_{1}$ such that
$\lim\limits_{n\rightarrow\infty}\|N_{{\bf f}_{n}}-\Phi\|_{E_{1}}=0.$
@@
$\Phi(z)=\lim\limits_{n\rightarrow\infty}N_{{\bf f}_{n}}(z),$
for each $z\in\Delta$. We will show that there is an $f_{\Phi}\in\mathcal{U}$ such that $N_{f_{\Phi}}=\Phi$. Without loss of generality, since the Pre-Schwarzian derivative is affine invariant, we assume that ${\bf f}_{n}\in\mathcal{S}$ for all $n$. Noting that $\{{\bf f}_{n}\}$ is a normal family, then $\{{\bf f}_{n}\}$ contains a subsequences (still denoted by $\{{\bf f}_{n}\}$) which is locally uniformly convergence in $\Delta$. We denote by $g$ the limit of the subsequences, i.e.,
$g(z)=\lim\limits_{n\rightarrow\infty}{\bf f}_{n}(z),$
@@
$N_{g}(z)=\lim\limits_{n\rightarrow\infty}N_{{\bf f}_{n}}(z).$
On the other hand, for $\Phi\in E_{1}$, there is a locally univalent function $f_{\Phi}$ in $\Delta$ such that
@@ 2.2. Quasiconformal mappings
Here and later, the notation ${\text{Area}}(E)$ denotes the Lebesgue measure of set $E$ in the complex plane. On the other hand, by Lemma 2.9, we have
@@ 3. Integral means spectrum functionals on Teichmüller spaces
We say $\mu$ and $\nu$ in $\mathcal{M}({\Delta}^{*})$ are asymptotically equivalent if there exists some $\widetilde{\nu}$ such that $\widetilde{\nu}$ and $\nu$ are equivalent and $\widetilde{\nu}(z)-\mu(z)\rightarrow 0$ as $|z|\rightarrow 1^{+}$. The asymptotic equivalence of $\mu$ will be denoted by $\mu]_{AT}$. The universal asymptotic Teichmüller space $AT$ is the set of all the asymptotic equivalence classes $\mu]_{AT}$ of elements $\mu$ in $\mathcal{M}({\Delta}^{*})$. The Teichmüller distance $d_{AT}([\mu]_{AT},[\nu]_{AT})$ of two points $\mu]_{AT}$, $\nu]_{AT}$ in $AT$ is defined as
@@
We see from [31] that $|\Phi(t)|\leq 8$. Noting that $f(z,0)=z$ so that $\Phi(0)=0$, then by using Schwarz’s lemma, we obtain that $|\Phi(t)|\leq 8t\|\mu_{\mathbf{f}}\|_{\infty}$ for $t\in\mathcal{D}_{\mathbf{f}}$. On the other hand, from $f(z,1)=\mathbf{f}(z)$, we have $|N_{\mathbf{f}}(z)|\rho_{f_{\mu}(\Delta)}^{-1}(z)\leq 8\|\mu_{\mathbf{f}}\|_{\infty}$ for all $z\in f_{\mu}(\Delta)$. It follows from the fact $\mathbf{f}=f_{\nu}\circ f_{\mu}^{-1}$ and (2.1) that
@@
In view of Propposition 3.2 and 3.3, we can identify the universal Teichmüller space with $T_{1}$ or $T_{2}$. Let $\mathcal{S}_{q}^{\infty}$ be the class of all functions $f\in\mathcal{S}_{q}$ that have a quasiconformal extension (still denoted by $f$) to $\widehat{\mathbb{C}}$ with $f(\infty)=\infty$. We set
@@
From $\mathbf{h}=g\circ f^{-1}$ and (2.1), we have
@@
where $z_{r}=re^{i\arg z}$, and the integral is taken on the radial path from $z_{r}$ to $z$.
On the other hand, since for all $|z|\in(r,1)$, $|P(z)(1-|z|^{2})|<\varepsilon$, then we have
@@
From [41, Proposition 8], we see that
@@ 5.1. A final main theorem
is holomorphic in $\Delta$. For fixed $r\in(0,1)$, by [37, Theorem 2.4.8], we have
@@
When $t_{1}>0$, let $\beta=\beta_{f_{\mu}}(t_{1})>0$. Let $\varepsilon<\beta$ be a positive number, which will be fixed later. Then we see from the definition of integral means spectrum of $f_{\mu}$ that there is a sequence $\{r_{n}\}_{n=1}^{\infty}$ with $r_{n}<1$ and $r_{n}\rightarrow 1$ as $n\rightarrow\infty$ and such that
@@
It is obvious that $\mathcal{E}_{n}\bigcup\mathcal{F}_{n}=[0,2\pi)$ and $\mathcal{E}_{n}\bigcap\mathcal{F}_{n}=\emptyset.$ We denote
@@
Hence, from [5] or [2], [42], we see that $\mathbf{h}$ is bounded univalent in $\Omega$ and admits a quasiconformal extension (still denoted by $\mathbf{h}$) to $\widehat{\mathbb{C}}$ when $\varepsilon$ small enough. Now, we let $\varepsilon$ be small enough so that $\mathbf{h}$ has a quasiconformal extension to $\widehat{\mathbb{C}}$. Then, let $\mathbf{F}(z)=\mathbf{h}\circ f_{\mu}(z),z\in\widehat{\mathbb{C}}$ and let $\mathbb{F}=\mathbf{F}|_{\Delta}$, we see that $\mathbb{F}$ belongs to $\mathcal{S}_{q}$, and for any $r\in(0,1)$, $\theta\in[0,2\pi)$, it holds that
$|\mathbb{F}^{\prime}(re^{i\theta})|^{t}=|\mathbf{h}^{\prime}\circ f_{\mu}(re^{i\theta})|^{t}|f^{\prime}_{\mu}(re^{i\theta})|^{t}=|f^{\prime}_{\mu}(re^{i\theta})|^{t+t\varepsilon}.$
@@
This corollary tells us that the extremal function for $B_{b}(t)$ can not be from the class $\mathcal{S}_{q}$ when $t\neq 0$. This partially answer part (3) of Question 1.13.
5.2. Remarks
@@
Corollary 5.5 tells us that, if the functional $I_{\overline{\mathcal{T}}}$ attains a maximum on $\overline{\mathcal{T}}$ for $t\neq 0$, then there is an extremal function $f$ for $B_{b}(t)$ whose Pre-Schwarzian derivative $N_{f}$ lies in $\partial\mathcal{T}$. In fact, for some special cases, we know that $I_{\overline{\mathcal{T}}}$ attains a maximum at certain point in $\partial{\mathcal{T}}$. For example,
@@
For each $t\in\mathbb{R}$, $B_{b}(t)$ has at least one extremal function whose Pre-Schwarzian derivative lies in $\partial\mathcal{T}$.
Remark 5.12.
From Proposition 4.7, we can identify the universal asymptotic Teichmüller space $AT$ with $\widetilde{\mathbf{N}}_{q}$. Based on Conjecture 5.11, we propose the following
Conjecture 5.13.
For each $t\in\mathbb{R}$, $B_{b}(t)$ has at least one extremal function $f$ such that $N_{f}]_{E_{1}}$ lies in the boundary of the universal asymptotic Teichmüller space $\widetilde{\mathbf{N}}_{q}$ in $E_{1}/E_{1,0}$.